#1 When a family with two adults and three children bought tickets for a movie, they paid a total of $27.25. The next family in line, with two children and three adults, paid $32.25 for the same movie.
a) Write a system of equations with two variables describing the situation above. Define your variables. For example x= …, y=… b) Find the adult and child prices of the movie tickets.
#2 a) Give an example of a linear inequality. b) Would the answer to your inequality be one point? Why or why not?

Answer by stanbon(75887)

When a family with two adults and three children bought tickets for a movie, they paid a total of $27.25. The next family in line, with two children and three adults, paid $32.25 for the same movie.
a) Write a system of equations with two variables describing the situation above. Define your variables. For example x= …, y=…
------
Equations:
2a + 3c = 27.25
3a + 2c = 32.25
----
Multiply thru 1st by 3
Multiply thru 2nd by 2
----
6a + 9c = 3*27.25
6a + 4c = 2*32.25
------
Subtract and solve for "c":
5c = 17.25
c = $3.45 (price of a child's ticket)
-----
Solve for "a":
2a + 3c = 27.25
2a + 3*3.45 = 27.25
2a = 16.9
a = $8.45 (price of an adult's ticket)
===================
Cheers,
Stan H.
===============
b) Find the adult and child prices of the movie tickets.
#2 a) Give an example of a linear inequality. b) Would the answer to your inequality be one point? Why or why not?